OpenStudy (anonymous):

solve logx +log (x-1)=log 4x for x.

OpenStudy (mathstudent55):

Remember that the sum of logs is the log of the product. Use that to simplify the left side, and write it as one single log. Then if the log of something equals the log of something else, then the two things are equal.

OpenStudy (anonymous):

would it be log 10 ^x log 10^x-1

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

Use the formulae... \[\log_{a}x + \log_{a}y = \log_{a}xy \] Then remove log of both sides, you will get a quadratic in 'x'... Hope this helps :-)

OpenStudy (anonymous):

start with \[\log(x(x-1))=\log(4x)\]

OpenStudy (anonymous):

since the log is a one to one function this tells you \[x(x-1)=4x\] and this is the quadratic equation you need to solve

OpenStudy (anonymous):

would it be x^2 =5x? @satellite73

OpenStudy (anonymous):

@SamuelAlden917 can you help me out ?

OpenStudy (anonymous):

or @mathstudent55

OpenStudy (anonymous):

i got x=o , x=5 is that correct?

OpenStudy (mathstudent55):